# Class Distance Weighted Cross Entropy Loss for Classification of Disease Severity

### v1

@@ Abstract
Assessing disease severity involving ordinal classes, where each class represents increasing levels of severity, benefit from loss functions that account for this ordinal structure. Traditional categorical loss functions, like Cross-Entropy (CE), often perform suboptimally in these scenarios. To address this, we propose a novel loss function, Class Distance Weighted Cross-Entropy (CDW-CE), which penalizes misclassifications more harshly when classes are farther apart. We evaluated CDW-CE on the Labeled Images for Ulcerative Colitis (LIMUC) dataset using various deep architectures. Its performance was compared against several categorical and ordinal loss functions. To analyze the quality of latent representations, we used t-distributed stochastic neighbor embedding (t-SNE) visualizations and quantified their clustering with the Silhouette Score. We also compared Class Activation Maps (CAM) generated by models trained with CDW-CE and CE loss, incorporating domain expert feedback to evaluate alignment with expert knowledge. Our results show that CDW-CE consistently improves performance in ordinal image classification tasks. It achieves higher Silhouette Scores, indicating better differentiation of class representations, and its CAM visualizations demonstrate a stronger focus on clinically significant regions, as confirmed by domain experts.
@@
ordinal classification; convolutional neural network; computer-aided diagnosis; loss function
Introduction
@@
In ordinal classification tasks, it is common to use the distance relation between different classes. This relation is evident in many examples, including disease severity classification tasks where classes can correspond to the severity of disease, and their ordered nature can be exploited with custom loss functions [1, 2, 3, 4], instead of employing categorical loss functions.
@@
In the medical domain, disease classes are assigned according to the symptom severity, thus implying an ordinal relation between classes[5]. Misclassification between further separated classes is more detrimental than the misclassification of relatively closer classes. For example, in the context of a medical diagnosis tool, erroneously categorizing a patient as healthy when they have a severe condition is considerably more problematic than misclassification between a mild or moderate condition. On the other hand, commonly used loss functions consider each class in isolation and do not account for any distance information between them, and both misclassification cases will have the same impact on the error. Consequently, both types of misclassifications are accorded equal weight in terms of error, disregarding the practical impact. To address this, a possible approach is mathematically representing the distance information in a loss function, thereby penalizing misclassifications between classes that are further apart more severely than those closer.
@@
In our earlier work, we investigated application of ordinal loss functions for ulcerative colitis severity estimation [6]. In this work, we extend the proposed loss function, Class Distance Weighted Cross Entropy (CDW-CE) and demonstrate its capabilities. In addition, we expand the CDW-CE by introducing an additive margin term into the loss function, analyzing class discrimination properties of models trained with different loss functions, and demonstrating its effectiveness with supplementary experiments comparing against various loss functions in the literature.
@@
Ulcerative Colitis is a disease for which several deep learning-based models have been proposed to classify disease severity. In this problem, disease severity is represented in increasing sequential order, and classification models are trained to predict the disease severity classes corresponding to input images correctly. Labeled Images for Ulcerative Colitis (LIMUC) [7] dataset is a publicly available dataset with ordinal classes and imbalanced class distributions. We have used this dataset to comparatively evaluate CDW-CE and other standard loss functions namely Cross-Entropy (CE), Mean Squared Error (MSE), CORN framework [2], cross-entropy with an ordinal loss term (CO2) [4], and ordinal entropy loss (HO2) [4] from the literature.
@@
We leveraged t-distributed stochastic neighbor embedding (t-SNE) plots to analyze the latent representations of models trained with CDW-CE and CE. We have also calculated the Silhouette Score of the t-SNE plots to assess the latent representation quality of different methods. Class Activation Map (CAM) outputs from CDW-CE and CE methods and raw endoscopic images are presented to medical experts, eliciting their opinions on the alignment of CAM outputs with observed symptoms.
The main contribution of this paper is the proposal of an advanced ordinal loss function, Class Distance Weighted Cross-Entropy (CDW-CE), with hyperparameters of $\alpha$, the class distance penalization power term, and margin, the separation robustness enforcer. As detailed in the experimental evaluation section, these hyperparameters can be tuned to achieve better results.
Related Work
Ordinal categories are frequently observed in real-world situations [8], particularly in the medical domain [9], such as disease severity classifications [10]. For ordinal classification problems, tailored methodologies, including novel loss functions that take ordinality into account, are reported to outperform traditional methods [11]. This section provides an overview of methods making use of such loss functions.
@@
An end-to-end deep learning method for ordinal regression, focusing on age estimation from facial images, was introduced in [12]. This approach involved transforming ordinal regression into a series of binary classification sub-problems and employing a multiple-output CNN for joint task resolution. By using $N-1$ output nodes for $N$ classes and applying label extension to the target class, their method is reported to perform better than other ordinal regression techniques, such as metric learning. However, this approach exhibited inconsistencies in output ranking within the subtasks. Consistent Rank Logits (CORAL) framework [1] addresses rank-inconsistencies in extended binary classification models. This framework tackles the problem of inconsistent classifier predictions in neural networks, where model confidence should ideally decrease as the rank increases. CORAL differs from the method proposed in [12], employing weight sharing in the penultimate layer during training, ensuring classifier consistency. The authors demonstrated the effectiveness of CORAL when integrated into common CNN architectures like ResNet, resulting in enhanced predictive performance for age estimation tasks. Conditional Ordinal Regression for Neural Network (CORN) framework [2] relaxes the constraints on the penultimate layer of the CORAL framework and incorporates conditional probabilities. Experiments on various datasets, including MORPH-2, AFAD, AES, and FIREMAN, showed that CORN outperformed previous methods. However, as mentioned earlier, these techniques need modification of both model architecture and labeling structure. An alternative approach for ordinal classification is integrating uni-modality distribution into the model’s output predictions, penalizing inconsistencies in posterior probability distribution between neighboring labels alongside the primary loss function, often using cross-entropy.
@@
Non-parametric ordinal loss [3] encourages output probabilities to follow an unimodal distribution by penalizing inconsistent probability pairs in the output. This approach was validated on various tasks, including breast cancer grading, predicting the decade of a photograph, and age estimation. An extension of this technique by adding unimodality losses to cross-entropy (CO2) and entropy losses (HO2) for the final loss function was proposed in [4], and it was validated using the Herlev dataset [13], consisting of cervical cell images, and various CNN architectures. Both studies reported improved performance compared to the method in [12] and cross-entropy loss.
CO2 and HO2 losses are frequently used for comparison in this study. In these equations, CE and H refer to Cross entropy and Entropy losses, respectively, $y_{n}$ is ground truth, $\hat{y}_{n}$ is model predictions, $K$ is the total number of classes, $\lambda$ is a parameter that determines the strength of the unimodal loss, and $\delta$ is the margin term.
@@
An alternative approach involves using regression to estimate a single continuous value at the output or applying a sigmoid activation function to constrain predictions within the range of [0, 1], which are then transformed into discrete levels using thresholds or probability distributions. Beckham et al. [14] proposed adding a single-node layer on top of the final layer with squared-error loss, suggesting that processing the final output through a sigmoid function followed by multiplication by (K - 1) yields better results, with rounding to the nearest integer for inference. They reported improved performance compared to standard cross-entropy loss. Polat et al. [15] developed a regression-based deep learning system, leveraging convolutional neural networks for enhanced accuracy and robustness in classifying Ulcerative Colitis severity. However, It is worth noting that regression-based approaches have generally been reported to exhibit inferior performance in various studies [4], [3].
Methods
@@
This work introduces an ordinal loss function designed to improve model performance in ordinal classification tasks, as previously discussed. This section provides detailed information on developing our custom loss function, the experimental setup, and the dataset used.
@@
The main idea behind employing an ordinal loss function is demonstrated in Figure 1. Unlike categorical loss functions, such as cross-entropy loss, which treat bimodal and unimodal distributions the same, an ordinal loss function can address issues specific to ordinal datasets. Mayo Endoscopic Scoring (MES) is used to classify the severity of Ulcerative Colitis, an inflammatory bowel disease. MES classes are ordinal, and thus, the misclassification of MES-2 (moderate disease severity) as MES-0 (no disease, healthy) is comparatively more harmful than the misclassification of adjacent classes (moderate and mild disease severities). We aim to encourage deep learning models to predict classes correctly while resembling an unimodal distribution to improve the robustness of the predictions.
@@
Class Distance Weighted Cross Entropy Loss
@@
Adding a margin to the loss function enforces intra-class compactness and increases inter-class distances, as shown in ArcFace [16], CosFace [17], SphereFace [18], and NormFace [19]. This is incorporated into CDW-CE loss by adding a margin to the class probabilities. CDW-CE with margin loss (Equation 4) is expected to improve the overall classification performance. However, this performance improvement comes at the cost of increased hyperparameter tuning and training time, as the margin term needs to be tuned for the specific dataset and task.
@@
Experimental Evaluation
@@
We used LIMUC [7] dataset and employed ResNet18 [20], Inception-v3 [21], and MobileNet-v3-large [22] architectures for experimental evaluation. In addition to the classification performance metrics, we have also calculated remission scores, a score used by medical experts that bins lower MES classes of 0 and 1 and higher MES classes of 2 and 3 together. This remission score is essential for domain experts aiming to classify whether a case is healthy or contains mild symptoms, which corresponds to the remission class, or contains moderate to severe symptoms, which is the non-remission class. We report the comparative performance evaluation of various methods using Quadratic Weighted Kappa (QWK) for multi-class classification, Kappa for binary (remission) classification, and accuracy, precision, recall, and F1 scores.
The proposed ordinal loss function Class Distance Weighted Cross Entropy (CDW-CE) evaluates the performance of a given prediction according to the ordinal nature and penalizes predictions further from the ground truth. The degree of penalization is adjustable through the $\alpha$ hyperparameter, which is a power term above the class distance coefficient. We analyzed the impact of $\alpha$ by varying it across integer values within the range of 1 to 10. For each model (ResNet18, Inception-v3, and MobileNet-v3-large), we found the best-performing power term values as 5, 6, and 7, respectively. These experiments are summarized in the Figure 2.
@@
We have implemented CDW-CE and other loss functions from the literature such as Cross-Entropy (CE), Mean Squared Error (MSE), CORN framework [2], cross-entropy with an ordinal loss term (CO2) [4], and ordinal entropy loss (HO2) [4]. We used Cross Entropy (CE) as our baseline and experimentally evaluated the CDW-CE against other loss functions.
@@
Dataset
We conducted our experiments on the LIMUC [7] dataset, which comprises endoscopic images with varying severity of Ulcerative Colitis (UC), ranked in increasing order from zero to three. UC is one of the two types of inflammatory bowel disease (IBD). Domain experts, i.e., medical doctors, classify whether the disease is present and rate its severity. The corresponding classes of healthy, mild, moderate, and severe disease conditions align with the ordinal levels of Mayo Endoscopic Scoring (MES) of zero to three. While there are several scoring systems for assessing UC from endoscopic images, the Mayo Endoscopic Score (MES) is the most widely used one [23]. MES assigns an ordinal label ranging from 0 (healthy instance) to 3 (severe symptoms) for symptom severity assessment of UC. Medical experts also employ a remission state scoring system, consisting of two classes, to distinguish between the presence (MES 2 and 3) and absence (MES 0 and 1) of UC.

### v2

@@ Abstract
Assessing disease severity with ordinal classes, where each class reflects increasing severity levels, benefits from loss functions designed for this ordinal structure. Traditional categorical loss functions, like Cross-Entropy (CE), often perform suboptimally in these scenarios. To address this, we propose a novel loss function, Class Distance Weighted Cross-Entropy (CDW-CE), which penalizes misclassifications more severely when the predicted and actual classes are farther apart. We evaluated CDW-CE using various deep architectures, comparing its performance against several categorical and ordinal loss functions. To assess the quality of latent representations, we used t-distributed stochastic neighbor embedding (t-SNE) and uniform manifold approximation and projection (UMAP) visualizations, quantified the clustering quality using the Silhouette Score, and compared Class Activation Maps (CAM) generated by models trained with CDW-CE and CE loss. Feedback from domain experts was incorporated to evaluate how well model attention aligns with expert opinion. Our results show that CDW-CE consistently improves performance in ordinal image classification tasks. It achieves higher Silhouette Scores, indicating better class discrimination capability, and its CAM visualizations show a stronger focus on clinically significant regions, as validated by domain experts. Receiver operator characteristics (ROC) curves and the area under the curve (AUC) scores highlight that CDW-CE outperforms other loss functions, including prominent ordinal loss functions from the literature.
@@
ordinal classification , convolutional neural network, computer-aided diagnosis, loss function
journal: Expert Systems With Applications
[…]
[inst1]organization=Graduate School of Informatics, addressline=Middle East Technical University, city=Ankara, country=Türkiye
1 Introduction
@@
In ordinal classification tasks, it is common to use the distance relation between different classes. This relation is evident in many examples, including disease severity classification tasks where classes can correspond to the severity of disease, and their ordered nature can be exploited with custom loss functions (Cao et al., 2020; Shi et al., 2023; Belharbi et al., 2019; Albuquerque et al., 2021), instead of employing categorical loss functions.
@@
In the medical domain, disease classes are assigned according to the symptom severity, thus implying an ordinal relation between classes (Kelley and Dozois, 2014). Misclassification between further separated classes is more detrimental than the misclassification of relatively closer classes. For example, in the context of a medical diagnosis tool, erroneously categorizing a patient as healthy when they have a severe condition is considerably more problematic than misclassification between a mild or moderate condition. On the other hand, commonly used loss functions consider each class in isolation and do not account for any distance information between them, and both misclassification cases will have the same impact on the error. Consequently, both types of misclassifications are accorded equal weight in terms of error, disregarding the practical impact. To address this, a possible approach is mathematically representing the distance information in a loss function, thereby penalizing misclassifications between classes that are further apart more severely than those closer.
@@
In our earlier work, we investigated application of ordinal loss functions for ulcerative colitis severity estimation (Polat et al., 2022a). In this work, we extend the proposed loss function, Class Distance Weighted Cross Entropy (CDW-CE) and demonstrate its capabilities. In addition, we expand the CDW-CE by introducing an additive margin term into the loss function, analyzing class discrimination properties of models trained with different loss functions, and demonstrating its effectiveness with supplementary experiments comparing against various loss functions in the literature.
@@
Ulcerative Colitis is a disease for which several deep learning-based models have been proposed to classify disease severity. In this problem, disease severity is represented in increasing sequential order, and classification models are trained to predict the disease severity classes corresponding to input images correctly. Labeled Images for Ulcerative Colitis (LIMUC) (Gorkem et al., 2022) dataset is a publicly available dataset with ordinal classes and imbalanced class distributions. We have used this dataset to comparatively evaluate CDW-CE and other standard loss functions namely Cross-Entropy (CE), Mean Squared Error (MSE), CORN framework (Shi et al., 2023), cross-entropy with an ordinal loss term (CO2) (Albuquerque et al., 2021), and ordinal entropy loss (HO2) (Albuquerque et al., 2021) from the literature.
@@
We used t-distributed stochastic neighbor embedding (t-SNE) and uniform manifold approximation and projection (UMAP) plots to analyze the latent representations of models trained with CE, CO2, HO2 and CDW-CE. To evaluate the quality of these representations, we calculated the silhouette score, a metric that quantifies inter-class compactness and intra-class separation (Malepathirana et al., 2022), based on the t-SNE and UMAP plots. Additionally, Class Activation Map (CAM) outputs from CDW-CE and CE methods, alongside with raw endoscopic images, were presented to medical experts to gather their feedback on the alignment of CAM outputs with observed symptoms. We also employed receiver operating characteristics (ROC) curves and area under curve (AUC) performance measures to highlight the importance of the loss functions on model performance.
The main contribution of this paper is the introduction of an advanced ordinal loss function, Class Distance Weighted Cross-Entropy (CDW-CE). The proposed CDW-CE loss function outperforms other loss functions in classification performance, feature extraction and explainability with default power and margin values.
2 Related Work
Deep learning-based disease severity classification is a widely researched topic and various methods have been proposed. Alyamani (2024) employs conventional convolutional neural networks (CNN) with categorical loss functions for the Ulcerative Colitis classification task. Barbero-Gómez et al. (2025) emphasizes the limited explainability of CNN methods and employs class activation maps to improved interpretability. de la Torre et al. (2018) explores the use case of quadratic weighted kappa (QWK) as a loss function for ordinal data.
Ordinal categories are frequently observed in real-world situations (Ünal et al., 2021), particularly in the medical domain (Seveso et al., 2020), such as disease severity classifications (Barash et al., 2021). For ordinal classification problems, tailored methodologies, including novel loss functions that take ordinality into account, are reported to outperform traditional methods (Tang et al., 2023). Ben-David (2008) emphasizes the importance of employing performance metrics, such as Cohen’s Kappa, for ordinal and multi-class classification problems. İsmail Kayadibi et al. (2023) propose using ROC curves and corresponding AUC scores to evaluate deep learning models trained on imbalanced datasets, which are common in the medical domain. Batchuluun et al. (2023) emphasize the use of class activation maps (CAMs) in deep learning models, particularly convolutional neural networks, to identify feature-level activations associated with the trained model’s attention. Similarly, Ahamed et al. (2024) employ CAM visualizations for better explainability, however, they do not employ medical domain expert feedback, leaving it for future studies.
This section provides an overview of methods that utilize such loss functions and evaluation methods.
@@ 2 Related Work
An end-to-end deep learning method for ordinal regression, focusing on age estimation from facial images, was introduced by Niu et al. (2016). This approach involved transforming ordinal regression into a series of binary classification sub-problems and employing a multiple-output CNN for joint task resolution. By using $N-1$ output nodes for $N$ classes and applying label extension to the target class, their method is reported to perform better than other ordinal regression techniques, such as metric learning. However, this approach exhibited inconsistencies in output ranking within the subtasks. Consistent Rank Logits (CORAL) framework (Cao et al., 2020) addresses rank-inconsistencies in extended binary classification models. This framework tackles the problem of inconsistent classifier predictions in neural networks, where model confidence should ideally decrease as the rank increases. CORAL differs from the method proposed by Niu et al. (2016), employing weight sharing in the penultimate layer during training, ensuring classifier consistency. The authors demonstrated the effectiveness of CORAL when integrated into common CNN architectures like ResNet, resulting in enhanced predictive performance for age estimation tasks. Conditional Ordinal Regression for Neural Networks (CORN) framework (Shi et al., 2023) relaxes the constraints on the penultimate layer of the CORAL framework and incorporates conditional probabilities. Experiments on various datasets, including MORPH-2, AFAD, AES, and FIREMAN, showed that CORN outperformed previous methods. However, as mentioned earlier, these techniques need modification of both model architecture and labeling structure. An alternative approach for ordinal classification is integrating uni-modality distribution into the model’s output predictions, penalizing inconsistencies in posterior probability distribution between neighboring labels alongside the primary loss function, often using cross-entropy.
@@
Non-parametric ordinal loss (Belharbi et al., 2019) encourages output probabilities to follow an unimodal distribution by penalizing inconsistent probability pairs in the output. This approach was validated on various tasks, including breast cancer grading, predicting the decade of a photograph, and age estimation. An extension of this technique by adding unimodality losses to cross-entropy (CO2) and entropy losses (HO2) for the final loss function was proposed by Albuquerque et al. (2021), and it was validated using the Herlev dataset (Jantzen and Dounias, 2006), consisting of cervical cell images, and various CNN architectures. Both studies reported improved performance compared to the method by Niu et al. (2016) and cross-entropy loss.
@@
An alternative approach involves using regression to estimate a single continuous value at the output or applying a sigmoid activation function to constrain predictions within the range of [0, 1], which are then transformed into discrete levels using thresholds or probability distributions. Beckham and Pal (2016) proposed adding a single-node layer on top of the final layer with squared-error loss, suggesting that processing the final output through a sigmoid function followed by multiplication by (K - 1) yields better results, with rounding to the nearest integer for inference. They reported improved performance compared to standard cross-entropy loss. Polat et al. (2022b) developed a regression-based deep learning system, leveraging convolutional neural networks for enhanced accuracy and robustness in classifying Ulcerative Colitis severity. However, It is worth noting that regression-based approaches have generally been reported to exhibit inferior performance in various studies (Albuquerque et al., 2021; Belharbi et al., 2019).
3 Methods
@@
This work introduces an ordinal loss function designed to improve model performance in ordinal classification tasks, as previously discussed. This section provides detailed information on developing our custom loss function, the experimental setup, evaluation metrics and the dataset used.
@@
The main idea behind employing an ordinal loss function is demonstrated in Figure 1. Unlike categorical loss functions, such as cross-entropy loss, which treat bimodal and unimodal distributions the same, an ordinal loss function can address issues specific to ordinal datasets. Mayo Endoscopic Scoring (MES) is used to classify the severity of Ulcerative Colitis, an inflammatory bowel disease. MES classes are ordinal, and thus, the misclassification of MES-3 (severe disease conditions) as MES-0 (no disease, healthy) is comparatively more harmful than the misclassification of adjacent classes (moderate and mild disease severities). We aim to encourage deep learning models to predict classes correctly while resembling an unimodal distribution to improve the robustness of the predictions.
@@
3.1 Class Distance Weighted Cross Entropy Loss
@@
Adding a margin to the loss function enforces intra-class compactness and increases inter-class distances, as shown in ArcFace (Deng et al., 2019), CosFace (Wang et al., 2018), SphereFace (Liu et al., 2017), and NormFace (Wang et al., 2017). This is incorporated into CDW-CE loss by adding a margin to the class probabilities. CDW-CE with margin loss (Equation 4) is expected to improve the overall classification performance.
@@
CO2 and HO2 losses, shown in Eqs. 5 and 6 respectively, are frequently used for comparison in this study. In these equations, CE and H refer to Cross entropy and Entropy losses, respectively, $y_{n}$ is ground truth, $\hat{y}_{n}$ is model predictions, $K$ is the total number of classes, $\lambda$ is a parameter that determines the strength of the unimodal loss, and $\delta$ is the margin term.
$\begin{split}\text{CO2}\left(y_{n},{\hat{y}}_{n}\right)=\text{CE}\left(y_{n},{\hat{y}}_{n}\right)+\lambda\sum_{k=0}^{K-1}{\mathbf{1}(k\geq k_{n}^{\ast})\text{RELU}(\delta+{\hat{y}}_{n\left(k+1\right)}-{\hat{y}}_{n(k)})}+\\ \lambda\sum_{k=0}^{K-1}{\mathbf{1}(k\leq k_{n}^{\ast})\text{RELU}(\delta+{\hat{y}}_{n\left(k\right)}-{\hat{y}}_{n(k+1)})}\end{split}$ (5)
$\begin{split}\text{HO2}\left(y_{n},{\hat{y}}_{n}\right)=\text{H}\left(y_{n},{\hat{y}}_{n}\right)+\lambda\sum_{k=0}^{K-1}{\mathbf{1}(k\geq k_{n}^{\ast})\text{RELU}(\delta+{\hat{y}}_{n\left(k+1\right)}-{\hat{y}}_{n(k)})}+\\ \lambda\sum_{k=0}^{K-1}{\mathbf{1}(k\leq k_{n}^{\ast})\text{RELU}(\delta+{\hat{y}}_{n\left(k\right)}-{\hat{y}}_{n(k+1)})}\end{split}$ (6)
By employing a power term for the distance of misclassification, the CDW-CE encourages the model to correctly form a unimodal prediction similar to CO2 and HO2. However, the power term penalizes further misclassifications more harshly thus preventing the model to favor majority-classes.
3.2 Evaluation Metrics
Evaluating the performance of deep learning methods is a challenging task that requires different approaches and solutions (Yilmaz and Demirhan, 2023). The classification performance of deep learning models was evaluated using accuracy, F1 score and quadratic weighted kappa (QWK). Additionally, ROC curves and their corresponding area under the curve(AUC) scores as well as confusion matrices were analyzed to compare different loss functions. Evaluation metrics used in this work are summarized with their associated descriptions and equations as well as widely used abbreviations in Table 1.
Metric Abbreviation Explanation
True
Positive
TP Correct prediction of a positive class
True
Negative
TN Correct prediction of a negative class
False
Positive
FP Incorrect prediction of a positive class
(type-1 error)
False
Negative
FN Incorrect prediction of a negative class
(type-2 error)
Accuracy Ratio of correct predictions to all samples
(TP+TN/TP+TN+FP+FN)
Precision Ratio of correct predicted positive samples to all positive predictions (TP/TP+FP)
Recall Ratio of correct predicted positive samples to actual positive samples (TP/TP+FN)
F1 Score Harmonic mean of precision and recall
(2*(Recall*Precision)///(Recall$+$Precision)
Kappa Ratio of observed accuracy to expected accuracy of random probability
Quadratic
Weighted
Kappa
QWK Kappa metric with more weight to further misclassifications between predicted and actual samples
Mean
Absolute
Error
MAE The average absolute difference between predicted and actual samples
Table 1: Evaluation metrics used.
@@ 3.2 Evaluation Metrics
In addition to the classification performance metrics, we have also calculated remission scores, a score used by medical experts that bins lower MES classes of 0 and 1 and higher MES classes of 2 and 3 together. This remission score is essential for domain experts aiming to classify whether a case is healthy or contains mild symptoms, which corresponds to the remission class, or contains moderate to severe symptoms, which is the non-remission class. We report the comparative performance evaluation of various methods using Quadratic Weighted Kappa (QWK) for multi-class classification, Kappa for binary (remission) classification, and accuracy, and F1 scores.
The classification task followed a one-vs-rest approach where the class with the highest probability was selected as the prediction. The deep learning models were also adapted for regression by modifying their final layers to have a single output node, without an activation function, and using the MSE loss function. This approach allows for the evaluation of regression-based models in this work. Mean absolute error (MAE), a typical baseline performance metric for regression tasks, was reported to compare regression based classification methods.
3.3 Experimental Setup
Figure 2: The flowchart of the experimental setup.
@@ 3.3 Experimental Setup
We used LIMUC (Gorkem et al., 2022) dataset and employed ResNet18 (He et al., 2016), Inception-v3 (Szegedy et al., 2015), and MobileNet-v3-large (Howard et al., 2019) architectures for experimental evaluation. We have implemented the proposed ordinal loss function Class Distance Weighted Cross Entropy (CDW-CE) and other loss functions from the literature such as Cross-Entropy (CE), Mean Squared Error (MSE), CORN framework (Shi et al., 2023), cross-entropy with an ordinal loss term (CO2) (Albuquerque et al., 2021), and ordinal entropy loss (HO2) (Albuquerque et al., 2021). We used Cross Entropy (CE) as our baseline and experimentally evaluated the CDW-CE against other loss functions. The flowchart of the experimental setup is provided in the Figure 2.
CDW-CE evaluates the performance of a given prediction according to the ordinal nature and penalizes predictions further from the ground truth. The degree of penalization is adjustable through the $\alpha$ hyperparameter, which is a power term above the class distance coefficient.
@@
The ability of a model to extract features correlates with how successfully it has been trained to understand and represent the underlying structure of the data. While feature extraction capability can be illustrated using t-SNE and UMAP plots, these plots do not allow direct numerical comparisons of embedding quality. To adress this, we used the Silhouette Score to quantitatively compare the feature extraction performance of the loss functions. For $N$ samples, the average Silhouette Score is calculated as shown in Eq. 7. where $a_{i}$ is the mean distance from the $i^{th}$ sample to other samples in the same cluster, and $b_{i}$ is the mean distance from the $i^{th}$ sample to samples in the nearest cluster.
$S=\frac{1}{N}\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{\max(a_{i},b_{i})}$ (7)
3.4 Dataset
We conducted our experiments on the LIMUC (Gorkem et al., 2022) dataset, which comprises endoscopic images with varying severity of Ulcerative Colitis (UC), ranked in increasing order from zero to three. UC is one of the two types of inflammatory bowel disease (IBD). Domain experts, i.e., medical doctors, classify whether the disease is present and rate its severity. The corresponding classes of healthy, mild, moderate, and severe disease conditions align with the ordinal levels of Mayo Endoscopic Scoring (MES) of zero to three. While there are several scoring systems for assessing UC from endoscopic images, the Mayo Endoscopic Score (MES) is the most widely used one (Satsangi et al., 2006). MES assigns an ordinal label ranging from 0 (healthy instance) to 3 (severe symptoms) for symptom severity assessment of UC. Medical experts also employ a remission state scoring system, consisting of two classes, to distinguish between the presence (MES 2 and 3) and absence (MES 0 and 1) of UC.